OpenStudy (eric_d):

A stone is projected vertically upward with a velocity of 10ms^-1 near the edge of the cliff.Given height of cliff=20mTaking g as 10ms^-2,find a)time taken for the stone to reach the ground b)velocity of stone when it hits the ground c)sketch the s-t,v-t and a-t graphs

OpenStudy (eric_d):

a)time taken for the stone to reach the ground

OpenStudy (eric_d):

@Abhisar

OpenStudy (eric_d):

I need to use s=ut+1/2at^2

OpenStudy (abhisar):

It's easy @eric_d ! Why don't you try first ?

OpenStudy (anonymous):

use t=2u/g

OpenStudy (eric_d):

I'm nt sure wth this Must i apply s=-20m

OpenStudy (abhisar):

First calculate the time taken to go to the maximum height and then back to the top of the cliff. Then calculate the time taken from the top of the cliff to the ground

OpenStudy (abhisar):



OpenStudy (abhisar):

Ball will go first A to B, calculate this time. Then ball will come back from B to C. It will take the same time. Now finally calculate time taken from C to D. Add all the time durations to get the final answer.

OpenStudy (abhisar):

getting my points ?

OpenStudy (eric_d):

s=-20m or +20m

OpenStudy (abhisar):

ok let's see. For A to B u=10, v=0, a=-10. using v=u+at, we get t= ?

OpenStudy (eric_d):

0=10+(-10)t t=1s

OpenStudy (abhisar):

Correct ! Now for B to C time taken will be same.

OpenStudy (abhisar):

also final velocity at C will also be equal to 10 m/s (by principal of conservation of mechanical energy)

OpenStudy (abhisar):

Now take u=10, s=20, a=10 and calculate s by s=ut+1/2at^2

OpenStudy (abhisar):

???

OpenStudy (eric_d):

20=10t+1/2(10)t^2

OpenStudy (abhisar):

yes, find t

OpenStudy (eric_d):